OpenStudy (anonymous):

convert 3pi/10 radians to degrees

OpenStudy (p0sitr0n):

p = 180 3pi/10 =x solve cross product

OpenStudy (p0sitr0n):

pi 10 = 180 /10 = 18 3 *18= 54